OpenStudy (anonymous):

Graph the function y = x^2 - 9. How many solutions are there? No Solution Infinite solutions Two Solutions One Solution

OpenStudy (anonymous):

What can x equal?

OpenStudy (tkhunny):

It's quadratic (Degree 2). Without any other information or restriction, there are 2.

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

i got no solutino

OpenStudy (solomonzelman):

Yes degree 2, with y-intercepts easy to calculate using `a²-b²=(a-b)(a+b)` You should listen to tkhunny, he is very knowledgeable !

OpenStudy (tkhunny):

If it is intended that we use the Real Numbers, and actually draw the graph on an x-y coordinate plane, then we have: x^2 - 9 = 0 ==> (x+3)(x-3) = 0 ==> x = 3 or x = -3 I count two solutions.

OpenStudy (anonymous):

y = 2x^2 + 1 what about this one?

OpenStudy (tkhunny):

Again, that depends on the rules that you haven't told us. Real Numbers? No solution. You can use the Quadratic Formula.

OpenStudy (solomonzelman):

imaginary solush :)